evaluate the following integral
∫ (2 cos 2x + sec^2 (x)) / sin 2x + tan x – 5 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{2cos2x+sec^{2}x}{sin2x+tanx-5}dx
sin2x+tanx-5 = t
(2cos2x+sec^{2}x)dx = dt
I = \int \frac{dt}{t}
I = log(t) + c
I = log(sin2x+tanx-5) + c